OpenStudy (anonymous):

Which of the following lines has a slope of -1/2?

OpenStudy (anonymous):

what are the options

OpenStudy (anonymous):

x + 2y = 0 x - 2y = 0 -x + 2y = 0

OpenStudy (anonymous):

you would first have to set each equation to y=mx+b. i believe it's the first one.

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

wecome!